# Celestial $Lw_{1+\infty}$ charges from a twistor action

### v1

@@ 3 Asymptotic twistor space and the Bramson-Tod integral formula
$\boxed{\mathcal{Q}_{s}=\int q^{s+2}\mathrm{d}q\wedge g|_{X_{(u,\lambda,\bar{\lambda})}}=\int q^{s+2}g_{1}\mathrm{d}q\mathrm{d}\bar{q}\,,}$ (3.21)
@@
$\begin{split}\partial_{u}\mathcal{Q}_{s}-\bar{\eth}\mathcal{Q}_{s-1}-(s+1)\bar{\sigma}\mathcal{Q}_{s-2}&=\int\left[q^{s+2}\partial_{u}g_{1}-q^{s+1}\bar{\eth}g_{1}-(s+1)\bar{\sigma}q^{s}g_{1}\right]\mathrm{d}q\mathrm{d}\bar{q}\\ &=\int q^{s+1}[q\partial_{u}g_{1}-\bar{\eth}g_{1}+\bar{\sigma}\partial_{q}g_{1}]\mathrm{d}q\mathrm{d}\bar{q}\,,\end{split}$ (3.24)
@@ 4 $Lw_{1+\infty}$ symmetry algebra
$\xi_{s,n}=\int q^{-n-1}\mathrm{d}q\wedge\delta h_{s}|_{X_{(u,\lambda,\bar{\lambda})}}\,,$ (4.9)
@@
The gauge parameter on the right-hand side is again of the form (4.13) with (4.10)-(4.14). We checked explicitly that the algebra closes for all $s$ at leading order in the fields. Furthermore, we also showed that the algebra closes up to $s=5$ to all orders in the fields. We refer to Appendix B for more details on these computations and explicit expressions for all $s$. We conclude that the symmetry algebra of the residual gauge diffeomorphisms $\xi$ is given by the $Lw_{1+\infty}$ algebra. Notice that, in this approach, there is no wedge condition on the parameters, which is in agreement with the phase space discussions on spacetime in [24, 25].
@@
$\delta_{s^{\prime}}\mathcal{Q}_{s}=\int q^{s+2}\mathrm{d}q\wedge\{g,\xi_{s^{\prime}}\}|_{X_{(u,\lambda,\bar{\lambda})}}\,,$ (4.19)
@@
$\begin{split}\delta_{s^{\prime}}\left(\partial_{u}\mathcal{Q}_{s}-\bar{\eth}\mathcal{Q}_{s-1}-(s+1)\bar{\sigma}\mathcal{Q}_{s-2}\right)&=\delta_{s^{\prime}}\int q^{s+1}\left[q\partial_{u}g_{1}-\bar{\eth}g_{1}+\bar{\sigma}\partial_{q}g_{1}\right]\mathrm{d}q\mathrm{d}\bar{q}\\ &=\frac{1}{\langle\lambda\hat{\lambda}\rangle}\delta_{s^{\prime}}\int q^{s+1}\left[\bar{\nabla}g_{\bar{\mathfrak{q}}}\right]_{\mathrm{D}\bar{\lambda}}\mathrm{d}q\mathrm{d}\bar{q}\,,\end{split}$ (4.24)
@@
$\delta_{s^{\prime}}\left(\partial_{u}\mathcal{Q}_{s}-\bar{\eth}\mathcal{Q}_{s-1}+(s+1)\bar{\sigma}\mathcal{Q}_{s-2}\right)=\frac{1}{\langle\hat{\lambda}\lambda\rangle^{2}}\int\partial_{\bar{q}}\left(q^{s+1}\{g_{0},\xi_{s^{\prime}}\}\right)\mathrm{d}q\mathrm{d}\bar{q}=0\,,$ (4.26)
@@ 5 Charge algebra
$\begin{split}\delta S=&\int_{\mathbb{PT}}\mathrm{D}^{3}Z\wedge\left[\left(\bar{\partial}h+\frac{1}{2}\{h,h\}\right)\wedge\delta g+\bar{\nabla}g\wedge\delta h\right]\\ &+\int_{\mathbb{PT}}\mathrm{D}^{3}Z\wedge\frac{\partial}{\partial\mu^{\dot{\alpha}}}\left(g\wedge\epsilon^{\dot{\alpha}\dot{\beta}}\frac{\partial h}{\partial\mu^{\dot{\beta}}}\wedge\delta h\right)+\int_{\mathbb{PT}}\mathrm{D}^{3}Z\wedge\bar{\partial}(g\wedge\delta h)\,.\end{split}$ (5.1)
The first term on the right-hand side corresponds to the equations of motion (2.18). The second term vanishes due to the gauge fixing condition (4.2). The last term is a total derivative term corresponding to the presymplectic potential. Taking one more variation of the latter, and integrating over the co-dimension $1$-surface $\Sigma=\{Z\in\mathbb{PT}|u_{B}-\bar{u}_{B}=\text{constant}\},$ [^5] The space of null twistors is the hypersurface $PN=\{Z\in\mathbb{PT}|u_{B}-\bar{u}_{B}=0\}$. we obtain the presymplectic form
$\boxed{\Omega(\delta g,\delta h)=\int_{\Sigma}\delta h\wedge\delta g\wedge D^{3}Z\,.}$ (5.2)
@@
$\delta_{\phi}\lrcorner\,\,\Omega+\delta H_{\phi}=0\quad,\quad H_{\phi}=\int_{\partial\Sigma}D^{3}Z\wedge h\phi\,.$ (5.3)
@@
where we have dropped the boundary terms associated with total $\partial_{q}$ terms because we assume the fibers have no boundary throughout. We also discarded a total $\partial_{u}$ term on the right-hand side, which reads explicitly as
$\int_{\Sigma}\mathrm{D}^{3}Z\wedge\partial_{u}\left(\partial_{q}\xi_{s}g\wedge\delta h\right)=\\ \left[\sum_{n=0}^{s}(n+1)\int_{\mathbb{CP}^{1}}\delta h_{0}\mathcal{Q}_{n-2}\left(\frac{u^{s-n}}{(s-n)!}\bar{\eth}^{s-n}\tau_{s}+\alpha_{n+1}[\bar{\sigma},\tau_{s}]\right)\mathrm{D}\lambda\mathrm{D}\bar{\lambda}\right]_{\iota^{0}}^{\iota^{+}}\,.$ (5.5)
This corner term vanishes for $s=0,1$ if we take the same falloffs in $u$ as in [77, 78]. Furthermore, this term vanishes for all $s$ if we restrict ourselves to the Schwartzian falloffs for the shear (4.28), as considered in [24, 25]. We proceed by neglecting this boundary term, but without assuming the wedge condition. We see from (5.4) that the charges associated with $\xi\in\Omega^{0,0}(\mathbb{PT},\mathcal{O}(2))$ are non-vanishing, hence confirming that the $Lw_{1+\infty}$ are indeed large gauge transformations on twistor space. In particular, the charge associated with $s=-1$ does not vanish, which means that the corresponding symmetry has a non-trivial action on phase space. The charges are non-integrable (i.e. non $\delta$-exact on the solution space) and can be split as
@@
$\boxed{H_{\xi}=\int_{\partial\Sigma}\xi g\wedge D^{3}Z,}\qquad\Xi_{\xi}[\delta]=\int_{\partial\Sigma}\delta\xi g\wedge D^{3}Z\,.$ (5.7)
The non-integrability of the charges, which appears for $s\geq 1$, can be traced back to the field dependence of the parameter $\xi_{s}$ due to the gauge condition (4.5). This might have been expected, as it is well known that the charges associated to the BMS symmetries in asymptotically flat spacetimes are already non-integrable [32, 33, 6, 34, 9, 12, 35, 10]. Several tools have been developed to deal with this non-integrability, including a modified bracket [6, 9, 41, 42, 43, 44, 45, 46] defined as
@@
where the bracket for vector fields was given in (4.15). Hence, the charges form a representation of the $Lw_{1+\infty}$ symmetries on twistor space. The consistency of this bracket with Jacobi identities follows directly from the consistency of the bracket (4.15). Notice that a different split between integrable and non-integrable pieces in (5.6)-(5.7),
@@
$\begin{split}\Omega&=\int_{\Sigma}\delta h_{0}\delta g\wedge\mathrm{d}q\wedge\mathrm{d}u\wedge\mathrm{D}\lambda\wedge\mathrm{D}\bar{\lambda}=-\int_{\mathscr{I}}\delta h_{0}\delta\Psi_{4}^{0}\mathrm{d}u\mathrm{D}\lambda\mathrm{D}\bar{\lambda}\,.\end{split}$ (5.12)
@@ 6 Discussion
Such symmetries are not expected to extend to the full Einstein vacuum equations in the bulk which are not believed to be integrable. There is therefore no direct evidence that these should be symmetries of the S-matrix for gravity save that they act naturally at the boundary. However, the integrability of the self-dual sector implies much more that symmetry of the S-matrix, rather instead the vanishing of the scattering of all positive helicity and the single negative and rest positive helicity amplitudes [55]. So, the lack of integrability does not preclude some natural action on the amplitude, although it is expected to be more complicated than the Ward identities arising from the extended BMS symmetries at low $s$.
@@ Acknowledgements
We would like to thank Laurent Freidel and Marc Geiller for insightful discussions on their work. We would also like to thank Tim Adamo, Giuseppe Bogna, and Simon Heuveline for helpful discussions. R.R. is also grateful to Glenn Barnich for related discussions during his Master’s thesis project in 2016. A.K. is supported by the STFC. R.R. is supported by the Titchmarsh Research Fellowship at the Mathematical Institute and by the Walker Early Career Fellowship in Mathematical Physics at Balliol College. A.Y.S. is supported by the STFC grant ST/X000761/1. This work is supported by the Simons Collaboration on Celestial Holography.

### v2

@@ 3 Asymptotic twistor space and the Bramson-Tod integral formula
$\boxed{\mathcal{Q}_{s}=\frac{1}{2\pi i}\int q^{s+2}\mathrm{d}q\wedge g|_{X_{(u,\lambda,\bar{\lambda})}}=\frac{1}{2\pi i}\int q^{s+2}g_{1}\mathrm{d}q\mathrm{d}\bar{q}\,,}$ (3.21)
@@
$\begin{split}\partial_{u}\mathcal{Q}_{s}-\bar{\eth}\mathcal{Q}_{s-1}-(s+1)\bar{\sigma}\mathcal{Q}_{s-2}&=\frac{1}{2\pi i}\int\left[q^{s+2}\partial_{u}g_{1}-q^{s+1}\bar{\eth}g_{1}-(s+1)\bar{\sigma}q^{s}g_{1}\right]\mathrm{d}q\mathrm{d}\bar{q}\\ &=\frac{1}{2\pi i}\int q^{s+1}[q\partial_{u}g_{1}-\bar{\eth}g_{1}+\bar{\sigma}\partial_{q}g_{1}]\mathrm{d}q\mathrm{d}\bar{q}\,,\end{split}$ (3.24)
@@
The integral formula (3.21) can be inverted to give an explicit form of $g$ which is only $\bar{\nabla}$-closed if the recursion relations (3.23) are satisfied. The Bramson-Tod representative is given by
$g=-\sum_{n=0}^{\infty}\partial_{u}\mathcal{Q}_{n-2}\frac{\bar{q}^{n}}{(1+q\bar{q})^{n+1}}\mathrm{D}\bar{\lambda}+\sum_{n=0}^{\infty}\mathcal{Q}_{n-2}\frac{(n+1)\bar{q}^{n}}{(1+q\bar{q})^{n+2}}\mathrm{d}\bar{q}\,.$ (3.36)
@@ 4 $Lw_{1+\infty}$ symmetry algebra
$\xi_{s,n}=\frac{1}{2\pi i}\int q^{-n-1}\mathrm{d}q\wedge\delta h_{s}|_{X_{(u,\lambda,\bar{\lambda})}}\,,$ (4.9)
@@
The gauge parameter on the right-hand side is again of the form (4.13) with (4.10)-(4.14). We checked explicitly that the algebra closes for all $s$ at leading order in the fields. Furthermore, we also showed that the algebra closes up to $s=5$ to all orders in the fields. We refer to Appendix B for more details on these computations and explicit expressions for all $s$. We conclude that the symmetry algebra of the residual gauge diffeomorphisms $\xi$ is given by the $Lw_{1+\infty}$ algebra. Notice that, in this approach, there is no wedge condition on the parameters, which is in agreement with the phase space discussions on spacetime in [24, 25]. [^5] We note that this algebra only becomes a Lie algebra if we restrict to the wedge as the Jacobi identity would not be satisfied. This is because $Lw_{1+\infty}$ has no Lie subalgebra keeping only $s\geq-1$. On spacetime phase-space one can also realize the generalized BMS algebra. We thank Nicolas Cresto and Laurent Freidel for pointing out this subtlety.
@@
$\delta_{s^{\prime}}\mathcal{Q}_{s}=\frac{1}{2\pi i}\int q^{s+2}\mathrm{d}q\wedge\{g,\xi_{s^{\prime}}\}|_{X_{(u,\lambda,\bar{\lambda})}}\,,$ (4.19)
@@
$\begin{split}\delta_{s^{\prime}}\left(\partial_{u}\mathcal{Q}_{s}-\bar{\eth}\mathcal{Q}_{s-1}-(s+1)\bar{\sigma}\mathcal{Q}_{s-2}\right)&=\frac{1}{2\pi i}\delta_{s^{\prime}}\int q^{s+1}\left[q\partial_{u}g_{1}-\bar{\eth}g_{1}+\bar{\sigma}\partial_{q}g_{1}\right]\mathrm{d}q\mathrm{d}\bar{q}\\ &=\frac{1}{2\pi i\langle\lambda\hat{\lambda}\rangle}\int q^{s+1}\delta_{s^{\prime}}\left[\bar{\nabla}g_{\bar{\mathfrak{q}}}\right]_{\mathrm{D}\bar{\lambda}}\mathrm{d}q\mathrm{d}\bar{q}\,,\end{split}$ (4.24)
@@
$\delta_{s^{\prime}}\left(\partial_{u}\mathcal{Q}_{s}-\bar{\eth}\mathcal{Q}_{s-1}-(s+1)\bar{\sigma}\mathcal{Q}_{s-2}\right)=\frac{1}{2\pi i\langle\hat{\lambda}\lambda\rangle^{2}}\int\partial_{\bar{q}}\left(q^{s+1}\{g_{0},\xi_{s^{\prime}}\}\right)\mathrm{d}q\mathrm{d}\bar{q}=0\,,$ (4.26)
@@ 5 Charge algebra
$\begin{split}\delta S=&\frac{1}{2\pi i}\int_{\mathbb{PT}}\mathrm{D}^{3}Z\wedge\left[\left(\bar{\partial}h+\frac{1}{2}\{h,h\}\right)\wedge\delta g+\bar{\nabla}g\wedge\delta h\right]\\ &+\frac{1}{2\pi i}\int_{\mathbb{PT}}\mathrm{D}^{3}Z\wedge\frac{\partial}{\partial\mu^{\dot{\alpha}}}\left(g\wedge\epsilon^{\dot{\alpha}\dot{\beta}}\frac{\partial h}{\partial\mu^{\dot{\beta}}}\wedge\delta h\right)+\frac{1}{2\pi i}\int_{\mathbb{PT}}\mathrm{D}^{3}Z\wedge\bar{\partial}(g\wedge\delta h)\,.\end{split}$ (5.1)
The first term on the right-hand side corresponds to the equations of motion (2.18). The second term vanishes due to the gauge fixing condition (4.2). The last term is a total derivative term from which the presymplectic potential can be extracted:
$\Theta[\delta]=\frac{1}{2\pi i}\int_{\Sigma}\theta[\delta]=\frac{1}{2\pi i}\int_{\Sigma}\delta h\wedge g\wedge D^{3}Z\,.$ (5.2)
Here the integration takes place over the co-dimension $1$-surface $\Sigma=\{Z\in\mathbb{PT}|u_{B}-\bar{u}_{B}=\text{constant}\}.$ [^6] The space of null twistors is the hypersurface $PN=\{Z\in\mathbb{PT}|u_{B}-\bar{u}_{B}=0\}$. Taking one more variation of $\Theta$, we obtain the presymplectic form
$\boxed{\Omega[\delta,\delta^{\prime}]=\frac{1}{2\pi i}\int_{\Sigma}\omega[\delta,\delta^{\prime}]=\frac{1}{2\pi i}\int_{\Sigma}\delta h\wedge\delta g\wedge D^{3}Z}$ (5.3)
with $\omega[\delta,\delta^{\prime}]=\delta^{\prime}\theta[\delta]-\delta\theta[\delta^{\prime}]$, the presymplectic current.
@@
$\delta_{\phi}\lrcorner\,\,\Omega+\delta F_{\phi}=0\,,\quad F_{\phi}=H_{\phi}|^{u_{B}=+\infty}_{u_{B}=-\infty}\,,\quad H_{\phi}=\frac{1}{2\pi i}\int_{\partial\Sigma}D^{3}Z\wedge h\phi\,.$ (5.4)
@@
$\delta_{\xi}\lrcorner\,\,\Omega+\delta F_{\xi}=0\,,\quad F_{\xi}=H_{\xi}|^{u_{B}=+\infty}_{u_{B}=-\infty}\,,\quad H_{\xi}=\frac{1}{2\pi i}\int_{\partial\Sigma}D^{3}Z\wedge\xi g\,,$ (5.5)
where we have dropped the boundary terms associated with total $\partial_{q}$ terms because we assume the fibers have no boundary throughout. There is also a total $\partial_{u}$ term, which read explicitly as
$\frac{1}{2\pi i}\int_{\Sigma}\mathrm{D}^{3}Z\wedge\partial_{u}\left(\partial_{q}\xi_{s}g\wedge\delta h\right)=\\ \left[\sum_{n=0}^{s}(n+1)\int_{\mathbb{CP}^{1}}\delta h_{0}\mathcal{Q}_{n-2}\left(\frac{u^{s-n}}{(s-n)!}\bar{\eth}^{s-n}\tau_{s}+\alpha_{n+1}[\bar{\sigma},\tau_{s}]\right)\mathrm{D}\lambda\mathrm{D}\bar{\lambda}\right]_{u_{B}=-\infty}^{u_{B}=+\infty}\,.$ (5.6)
This corner term vanishes for $s=0,1$ if we take the same falloffs in $u$ as in [77, 78]. Furthermore, this term vanishes for all $s$ if we restrict ourselves to the Schwartzian falloffs for the shear (4.28), as considered in [24, 25]. We proceed by neglecting this boundary term, but without necessarily assuming the wedge condition.
We see from (5.6) that the charges associated with $\xi\in\Omega^{0,0}(\mathbb{PT},\mathcal{O}(2))$ are non-vanishing, hence confirming that the $Lw_{1+\infty}$ are indeed large gauge transformations on twistor space. In particular, the charge associated with $s=-1$ does not vanish, which means that the corresponding symmetry has a non-trivial action on phase space. Furthermore, one can easily show that the charges form a representation of the $Lw_{1+\infty}$ symmetries, i.e.
$\boxed{\delta_{\xi}F_{\xi^{\prime}}=F_{\{\xi,\xi^{\prime}\}_{*}}}$ (5.7)
where the bracket for the residual gauge parameters on the right-hand side was defined in (4.15). This important result could only be obtained at linear order from a spacetime approach [24, 25], demonstrating the power of the twistor methods presented here. Notice that restricting (5.7) to $s=0,1$ reproduces the BMS flux algebra, in agreement with [77, 78].
It is also instructive to derive the surface charge expressions at finite value of $u_{B}$, which will coincide with the surface charges integrated on a finite cut of $\mathscr{I}$ after translation to spacetime, see next paragraph. We have [29, 30, 31]
@@
where $k_{\xi}$ is a co-dimension $2$ form on twistor space, which can be integrated over $\partial\Sigma=\{Z\in\mathbb{PT}|u_{B}-\bar{u}_{B}=0,u_{B}=\text{constant}\}$. The resulting charges are non-integrable (i.e. non $\delta$-exact on the solution space) and can be split as
@@
$\boxed{H_{\xi}=\frac{1}{2\pi i}\int_{\partial\Sigma}D^{3}Z\wedge\xi g}\,,\qquad\Xi_{\xi}[\delta]=H_{\delta\xi}+\frac{1}{2\pi i}\int_{\partial\Sigma}\underline{\xi}\lrcorner\,\theta[\delta]\,.$ (5.10)
where $\underline{\xi}=\{\xi,\,\cdot\,\}$ is the Hamiltonian vector field generated by $\xi$. The explicit expression for the second term in $\Xi_{\xi}[\delta]$ can be read of from (5.6). The integrable/non-integrable split adopted here corresponds to the Noetherian split discussed e.g. in [45]. The non-integrable piece $\Xi_{\xi}$ vanishes when $u\to\pm\infty$ for $s=0,1$ if we take the same falloffs in $u$ as in [77, 78], and vanishes for all $s$ if we restrict ourselves to the Schwartzian falloffs for the shear (4.28), as considered in [24, 25]. At finite $u$, the non-integrability of the charges appears for $s\geq 1$. This might have been expected, as it is well known that the charges associated to the BMS symmetries in asymptotically flat spacetimes are already non-integrable [32, 33, 6, 34, 9, 12, 35, 10]. Several tools have been developed to deal with this non-integrability, including a modified bracket [6, 9, 41, 42, 43, 44, 45, 46] defined as
@@
where the bracket for vector fields was given in (4.15). Hence, the charges form a representation of the $Lw_{1+\infty}$ symmetries on twistor space. The consistency of this bracket with Jacobi identities follows directly from the consistency of the bracket (4.15). Using the above arguments, one can check that the result (5.12) consistently reduces to (5.7) when taking $u\to\pm\infty$ and assuming the appropriate falloff conditions in $u$. Finally, notice that a different split between integrable and non-integrable pieces in (5.9)-(5.10),
@@
$\begin{split}\Omega&=\frac{1}{2\pi i}\int_{\Sigma}\delta h_{0}\delta g\wedge\mathrm{d}q\wedge\mathrm{d}u\wedge\mathrm{D}\lambda\wedge\mathrm{D}\bar{\lambda}=-\int_{\mathscr{I}}\delta h_{0}\delta\Psi_{4}^{0}\mathrm{d}u\mathrm{D}\lambda\mathrm{D}\bar{\lambda}\,.\end{split}$ (5.15)
@@ 6 Discussion
Such symmetries are not expected to extend to the full Einstein vacuum equations in the bulk which are not believed to be integrable. There is therefore no direct evidence that these should be symmetries of the S-matrix for gravity save that they act naturally at the boundary. Indeed, the integrability of the self-dual sector implies much more than symmetry of the S-matrix, it implies the vanishing of all scattering amplitudes except for the 3-point $\overline{\text{MHV}}$ amplitude and the 1-loop all positive external helicity amplitude [55]. So, the lack of integrability does not preclude some natural action on the amplitude, although it is expected to be more complicated than the Ward identities arising from the extended BMS symmetries at low $s$.
@@ Acknowledgements
We would like to thank Nicolas Cresto, Laurent Freidel and Marc Geiller for insightful discussions on their work. We would also like to thank Tim Adamo, Giuseppe Bogna, and Simon Heuveline for helpful discussions. R.R. is also grateful to Glenn Barnich for related discussions during his Master’s thesis project in 2016. A.K. is supported by the STFC. R.R. is supported by the Titchmarsh Research Fellowship at the Mathematical Institute and by the Walker Early Career Fellowship in Mathematical Physics at Balliol College. A.Y.S. is supported by the STFC grant ST/X000761/1. This work is supported by the Simons Collaboration on Celestial Holography.